Circular No. 8023

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2002jh Further to IAUC 8006, W. M. Wood-Vasey, G. Aldering, P. Nugent,

and K. Li report the discovery of an apparent supernova (mag 18.4) on unfiltered NEAT images taken with the Palomar 1.2-m Schmidt telescope on Nov. 12.48, 12.49, and 12.52 UT. Confirming NEAT images were made on Nov. 15 (mag 18.3), 17 (mag 18.5), and 25 (mag 18.6). The new object is located at α = 4^{h}27^{m}29^{s}.57, δ = +6^{o}01′34′′.2 (equinox 2000.0) which is 5′′ east and 3′′ south of the center of the apparent host galaxy. No point source is visible at this location in previous NEAT images taken on Oct. 31 (limiting mag 20.4 at S/N = 5).

SUPERNOVAE 2002je AND 2002jg T. Matheson, P. Challis, and R. Kirshner, Harvard-Smithsonian Center

for Astrophysics, report that on spectrograms (range 370–750 nm) obtained of SN 2002je (cf. IAUC 8021) and 2002jg (cf. IAUC 8022), obtained by P. Berlind on Nov. 28.13 and 28.08 UT with the Mt. Hopkins 1.5-m telescope (+ FAST spectrograph). SN 2002je is a type-II supernova near maximum light; the spectrum consists of a continuum with superposed P-Cyg lines of hydrogen. Narrow Hα emission superposed on the spectrum of SN 2002je indicates a recession velocity of 7470 km/s for the host galaxy; the expansion velocity derived from the minimum of the Hβ line is ∼ 11300 km/s. SN 2002jg is a reddened type-Ia supernova near maximum; adopting the NASA/IPAC Extragalactic Database recession velocity of 4718 km/s for the host galaxy, the supernova expansion velocity is ∼ 11100 km/s for Si II (rest 635.5 nm). The spectral-feature age of SN 2002jg is 1 ± 2 days before maximum light. The spectrum shows a relatively strong Na I D absorption at the velocity of of the host galaxy (equivalent width 0.2 nm). The dust maps of Schlegel et al. (1998, Ap.J. 500, 525) indicate a Galactic reddening of E(B − V) = 0.66. In addition, the overall shape of the spectrum is flat toward the blue, implying that SN 2002jg is extinguished by dust.

SATELLITES OF JUPITER Corrigendum. On IAUC 7998, the correct spelling of the name ap-

proved by the IAU Working Group on Planetary System Nomenclature for Jupiter XIX is ‘Megaclite’.

2002 November 29  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Daniel W. E. Green